OpenStudy (anonymous):

PLEASE, can you help me with this? I have no idea what to do. Which of the following points lie in the solution set to the following system of inequalities? y is less then or equal to x - 5 y is greater then or equal to -x - 4 1. (-5, 2) 2. (5, -2) 3. (-5, -2) 4. (5, 2)

OpenStudy (anonymous):

Could you clarify the inequalities? Right now there aren't any symbols so it's hard to tell which ordered pair will be a solution.

OpenStudy (anonymous):

Woops. Sorry. Didn't realize there weren't any there.

OpenStudy (anonymous):

let's start with the first one: \[y \le x-5\] how i would go about this problem is just to plug in the answers, and eliminate the wrong choices. i'll try answer 2 first (5, -2) that would be \[-2 \le 5-5\] or just \[-2 \le 0\] is that true? yes, which means this could be our answer choice. let's try it with the second equation now \[y \ge -x-4\] so just like the first problem, we'll plug in (5, -2) \[-2\ge -5-4\] or just \[-2 \ge -9\] which is true, so that means that answer choice 3. (5, -2) is your answer.